### v2

# The Generalized Dirac Equation in the Metric-Affine Spacetimes

Abstract
We discuss the most general form of Dirac equation in the non-Riemannian spacetimes containing curvature, torsion and non-metricity. It includes all bases of the Clifford algebra $cl(1,3)$ within the spinor connection. We adopt two approaches. First, the generalized Dirac equation is directly formulated by applying the minimal coupling prescription to the original Dirac equation. It is referred to as the direct Dirac equation for seek of clarity and to preserve the tractability. Second, through the application of variational calculation to the original Dirac Lagrangian, the resulting Dirac equation is referred to as the variational Dirac equation. A consistency crosscheck is performed between these two approaches, leading to novel constraints on the arbitrary coupling constants appearing in the covariant derivative of spinor. Following short analysis on the generalized Dirac Lagrangian, it is observed that two of the novel terms give rise to a shift in the spinor mass by sensing its handedness.
Keywords: Curvature, torsion, non-metricity, metric-affine spacetimes, Dirac equation, spinor bundle, spinor connection.
1 Introduction
The standard half-integral spin particle Dirac equation is foundational for quantum field studies. It was originally formulated in flat Minkowski spacetime by combining quantum mechanics and special relativity [1, 2, 3, 4]. Spacetime symmetric potentials on the other hand, require representation of the Dirac equation in the curved spacetime, where the partial derivatives are replaced with covariant derivatives that include the Christoffel symbols compatible with local Lorentz symmetry. This substitution, often called the minimal coupling prescription, is essential for adapting the free Dirac Lagrangian to a generally covariant setting, ensuring invariance under local Lorentz transformations and coordinate diffeomorphisms [5, 6, 7, 8]. In terms of the exterior algebra, this amounts to the formal replacement $d\rightarrow D$ at the level of the Dirac operator, leading to a consistent coupling of spinor fields to the background gravitational field encoded in the orthonormal coframe and the spin connection 1-form [9]. In the standard Einstein-Dirac framework, the minimal coupling allows for a well-posed first-order field equation for fermions on a torsion-free, metric-compatible manifold and ensures that in locally inertial frames the curved-spacetime Dirac equation reduces to its flat-space form, preserving the equivalence principle at the quantum level.
On the other hand, the various constraints on the metric and the connection (metric compatibility and zero-torsion as in General Relativity), which are defined independently of each other, do not arise from any fundamental principle. They are a priori constraints corresponding to the assumptions and postulates present in the theory describing a physical phenomenon. Therefore, there are researches in the non-Riemannian spacetimes. Consider the Einstein-Cartan theory. It refines general relativity by permitting a non-symmetric affine connection, whose antisymmetric part, the torsion tensor, is algebraically sourced by the intrinsic spin density of matter [10, 11]. This coupling between torsion and spin modifies both the gravitational field equations and the Dirac equation itself, yielding nonlinear corrections to the latter. The resulting Dirac equation incorporates spintorsion coupling terms, which are cubic in spinor fields and act as effective self-interactions [12, 13]. Yet, nontrivial consistency problems persist in coupling spinor fields to spacetimes with torsion. The primary one is related to the minimal coupling prescription. Its naive application in the presence of torsion may results in non-conserved spin currents or ambiguities in defining gauge-invariant currents. This inconsistency problem simply arises due to a direct substitution $d\rightarrow D$ at the level of field equations. It fails to preserve the underlying symmetries and conservation laws [14, 15]. Minimal coupling must be implemented at the action level rather than in the equation of motion [16, 17, 18]. This guarantees that the geometrically induced torsion terms do not conflict with the gauge symmetries or physical conservation laws of the theory.
Another complication arises when the background geometry includes non-metricity, i.e., non-vanishing covariant derivative of the metric. In this case, the usual algebra of the gamma matrix and spinor structure may be ill-defined under parallel transport [19, 20, 21]. The definition of the Dirac operator becomes inherently ambiguous. Several methods in response to this difficulty, have been proposed. Some even succeed in partially restoring a form of minimal coupling. Among these, one finds the use of anholonomic frames, band-spinor representations of $GL(n,\mathbb{R})$, non-linear connection structures, world spinors, etc. [19, 20, 21, 22, 23, 24, 25, 26]. However, a fully consistent and unambiguous formulation in non-metric spacetimes remains a nontrivial problem. In the present paper, possibly the most general form of the Dirac equation at the background of the metric-affine spacetimes including non-metricity, torsion and curvature is considered. A new resolution is introduced. The Dirac equation accordingly is derived from the first principles.
The paper is organized as follows. In Section 2, mathematical tools required for the metric-affine geometries in terms of the exterior algebra. Section 3 reviews the Clifford algebra and the spinor bundle. The evolution of the Dirac equation from Dirac’s original work to the Einstein-Cartan-Dirac theory is revisited by following the minimal coupling strategy in Section 4. The inconsistency problem in the Dirac equation is identified in this section. By a novel definition of covariant exterior derivative of spinor, $\mathbb{D}\psi$, Section 5 is devoted to presenting our solution. Note that, some constraints are applied to the coupling constants appearing in the definition of $\mathbb{D}\psi$. Two of our novel contributions in $\mathbb{D}\psi$ cause a shift in the spinor mass. Moreover, geometry may lead to mass differences between the spinor’s left- and right-handed components. Finally, Section 6 offers a detailed examination on this subject.
2 Metric-affine geometry
The triple $\{M,g,\nabla\}$ or $\{Q_{ab},T^{a},R^{a}{}_{b}\}$ defines a metric-affine geometry where $M$ is a four-dimensional orientable and differentiable manifold, $g$ is a non-degenerate symmetric metric, $\nabla$ is a full (or affine) connection or covariant derivative, $Q_{ab}$ is non-metricity 1-form, $T^{a}$ is torsion 2-form and $R^{a}{}_{b}$ is curvature 2-form [27, 28]. We denote the anholonomic metric-orthonormal coframe by $e^{a}$, then write the metric as $g=\eta_{ab}e^{a}\otimes e^{b}$ where $\eta_{ab}$ is the Minkowski metric with the signature (,+,+,+)(-,+,+,+). Latin indices are the anholonomic orthonormal (Lorentz) spacetime indices; $a,b,c,\cdots=0,1,2,3$. The affine connection is determined by the full connection 1-form $\omega^{a}{}_{b}$ by the definition $\nabla e^{a}:=-\omega^{a}{}_{b}\wedge e^{b}$ or $\nabla X_{a}:=X_{b}\otimes\omega^{b}{}_{a}$ where $\otimes$ and $\wedge$ denote the tensor product and exterior product, respectively, and $X_{a}$ is the orthonormal frame such that $e^{a}(X_{b})=\delta^{a}_{b}$ is the duality relation. Besides, $e^{a}$ is called the metric orthonormal (or shortly orthonormal) basis 1-form and the Cartan structure equations are given by tensor-valued non-metricity 1-form [^1] Here $\omega_{(ab)}=Q_{ab}$ is the reason why we adopt the factor $\frac{1}{2}$ in the definition of non-metricity, tensor-valued torsion 2-form and tensor-valued full (or non-Riemannian) curvature 2-form, respectively,
$Q_{ab}$ $:=-\frac{1}{2}D\eta_{ab}=-\frac{1}{2}\left(d\eta_{ab}-\omega^{c}{}_{a}\eta_{cb}-\omega^{c}{}_{b}\eta_{ac}\right)=\omega_{(ab)},$ (1a)
$T^{a}$ $:=De^{a}=de^{a}+\omega^{a}{}_{b}\wedge e^{b},$ (1b)
$R^{a}{}_{b}$ $:=D\omega^{a}{}_{b}:=d\omega^{a}{}_{b}+\omega^{a}{}_{c}\wedge\omega^{c}{}_{b}$, (1c)
where $d$ is the exterior derivative and $D$ is the covariant exterior derivative. We used the result $d\eta_{ab}=0$ in the orthonormal frame given by Eq.(1a). We utilize the usual notation: $(\bullet)$ and $\bullet$] mean to calculate the symmetric and anti-symmetric parts, respectively, with respect to the indices enclosed, i.e., $\omega_{(ab)}=\frac{1}{2}(\omega_{ab}+\omega_{ba})$ and $\omega_{[ab]}=\frac{1}{2}(\omega_{ab}-\omega_{ba})$. They satisfy the Bianchi identities
$DQ_{ab}=R_{(ab)},\qquad DT^{a}=R^{a}{}_{b}\wedge e^{b},\qquad DR^{a}{}_{b}=0.$ (2)
If need, one can pass to tensor notation easily by expanding the concerned exterior forms in components as follows
$\omega^{a}{}_{b}=\omega^{a}{}_{bc}e^{c},\qquad Q_{ab}=Q_{abc}e^{c},\qquad T^{a}=\frac{1}{2}T^{a}{}_{bc}e^{b}\wedge e^{c},\qquad R^{a}{}_{b}=\frac{1}{2}R^{a}{}_{bcd}e^{c}\wedge e^{d}.$ (3)
Thus, we deduce that $Q_{abc}=Q_{(ab)c}$ is symmetric at the first two indices, $T^{a}{}_{bc}=T^{a}{}_{[bc]}$ is anti-symmetric at the last two indices and $R_{abcd}=R_{ab[cd]}$ is anti-symmetric at the last two indices and asymmetric $R_{abcd}=R_{[ab]cd}+R_{(ab)cd}$ at the first two indices by definition. In the subsequent sections, the definitions of the first kind trace 1-form $Q:=\eta_{ab}Q^{ab}=Q^{a}{}_{ab}e^{b}$ and the second kind trace 1-form $P:=(\iota_{a}Q^{ab})e_{b}=Q^{a}{}_{ba}e^{b}$ of the non-metricity tensor, as well as the trace 1-form of the torsion $T:=\iota_{a}T^{a}$ will be frequently used where $\iota_{a}\equiv\iota_{X_{a}}$ denotes the interior derivative, $\iota_{b}e^{a}=\delta^{a}_{b}$. Furthermore, we would like to explicitly remind that $d\eta_{ab}=0$, $d\eta^{ab}=0$ and $d\delta^{a}_{b}=0$, but $D\eta_{ab}=-2Q_{ab}$, $D\eta^{ab}=+2Q^{ab}$ and $D\delta^{a}_{b}=0$. Accordingly, while it is free to move an index vertically in front of $d$ operator, one must pay special attention to raising or lowering an index in front of $D$ operator in non-metric geometries. Furthermore, the following identities are very useful for the calculations.
$D*e_{a}$ $=-Q\wedge*e_{a}+T^{b}\wedge*e_{ab}$ (4a)
$D*e_{ab}$ $=-Q\wedge*e_{ab}+T^{c}\wedge*e_{abc}$ (4b)
$D*e_{abc}$ $=-Q\wedge*e_{abc}+T^{d}\wedge*e_{abcd}$ (4c)
$D*e_{abcd}$ $=-Q\wedge*e_{abcd}$ (4d)
Nevertheless, the full connection 1-form, $\omega^{a}{}_{b}$, can be decomposed uniquely to a Riemannian piece, $\widetilde{\omega}_{ab}(g)$ determined by metric plus a non-Riemannian piece, $\mathrm{L}_{ab}(T,Q)$ determined by torsion and non-metricity [26, 29, 30]
$\omega_{ab}=\widetilde{\omega}_{ab}+\mathrm{L}_{ab},$ (5)
where $\widetilde{\omega}_{ab}=-\widetilde{\omega}_{ba}$ is the Levi-Civita connection 1-form,
$\widetilde{\omega}_{ab}=\frac{1}{2}\left[-\iota_{a}de_{b}+\iota_{b}de_{a}+(\iota_{a}\iota_{b}de_{c})e^{c}\right]\qquad\text{or}\qquad\widetilde{\omega}^{a}{}_{b}\wedge e^{b}=-de^{a}$ (6)
and $\mathrm{L}_{ab}$ is the tensor-valued defect (or distortion) 1-form,
$\mathrm{L}_{ab}:=\underbrace{\underbrace{\frac{1}{2}\left[\iota_{a}T_{b}-\iota_{b}T_{a}-(\iota_{a}\iota_{b}T_{c})e^{c}\right]}_{contortion}+\underbrace{(\imath_{b}Q_{ac}-\imath_{a}Q_{bc})e^{c}+Q_{ab}}_{disformation}}_{distortion}.$ (7)
In the literature it is common to define the tensor-valued contortion 1-form, $K_{ab}=-K_{ba}$, in terms of torsion 2-form
$K_{ab}=\frac{1}{2}\left[\iota_{a}T_{b}-\iota_{b}T_{a}-(\iota_{a}\iota_{b}T_{c})e^{c}\right]\qquad\text{or}\qquad K^{a}{}_{b}\wedge e^{b}=T^{a}.$ (8)
It is worthy to notice that the symmetric part of the affine connection is determined by only non-metricity, $\omega_{(ab)}=Q_{ab}$, the remainder of $\omega_{ab}$ is the anti-symmetric, $\omega_{[ab]}=\widetilde{\omega}_{ab}+K_{ab}+(\imath_{b}Q_{ac}-\imath_{a}Q_{bc})e^{c}$.
When non-metricity is set to zero, the full connection is called metric compatible in which case the symmetric parts of full connection and full curvature vanish. If torsion is also reset along with non-metricity, the full connection is named the Levi-Civita (or Riemannian) connection. An affine geometry is classified whether non-metricity, torsion and/or full curvature vanish or not [29].
3 Spinor bundle
The Clifford algebra (in fact the so-called Dirac algebra) $cl(1,3)$ is generated by the Minkowski vector space with signature (,+,+,+)(-,+,+,+). The set $\{I,\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3}\}$ of $4\times 4$ complex matrices could be the generators of $cl(1,3)$ which satisfy the condition
$\gamma_{a}\gamma_{b}+\gamma_{b}\gamma_{a}=2\eta_{ab}I.$ (9)
Besides, we represent the basis of $cl(1,3)$ by $\{I,\gamma_{a},\sigma_{ab},\gamma_{a}\gamma_{5},\gamma_{5}\}$ where $\sigma_{ab}:=(\gamma_{a}\gamma_{b}-\gamma_{b}\gamma_{a})/4$ and $\gamma_{5}:=\gamma_{0}\gamma_{1}\gamma_{2}\gamma_{3}$. We adhere the following representations of Dirac matrices
$\gamma_{0}=\begin{pmatrix}-iI_{2}&0\\ 0&iI_{2}\end{pmatrix},\qquad\gamma_{k}=\begin{pmatrix}0&i\sigma^{k}\\ -i\sigma^{k}&0\end{pmatrix}$, (10)
where $i=\sqrt{-1}$ is the imaginary unit, $I_{2}$ is the two-dimensional unit matrix, $\sigma^{k}$, $k=1,2,3$, are the Pauli matrices. There are relations between our matrices and Bjorken&\&Drell matrices [2] as $i\gamma_{0}=\beta$ and $\gamma_{0}\gamma_{k}=\alpha_{k}$. The bases of $cl(1,3)$ satify the following useful properties
$\gamma_{0}I^{\dagger}\gamma_{0}=-I,\qquad\gamma_{0}\gamma_{a}^{\dagger}\gamma_{0}=\gamma_{a},\qquad\gamma_{0}\sigma_{ab}^{\dagger}\gamma_{0}=\sigma_{ab},$
$\gamma_{0}(\gamma_{a}\gamma_{5})^{\dagger}\gamma_{0}=-\gamma_{a}\gamma_{5},\qquad\gamma_{0}\gamma_{5}^{\dagger}\gamma_{0}=-\gamma_{5},$ (11)
where denotes Hermitian conjugate. The commutation relations of the basis elements of $cl(1,3)$ are computed as
$[I,I]$ $=[I,\gamma_{a}]=[I,\sigma_{ab}]=[I,\gamma_{a}\gamma_{5}]=[I,\gamma_{5}]=0,$ (12a)
$[\gamma_{a},\gamma_{b}]$ $=4\sigma_{ab},$ (12b)
$[\gamma_{a},\sigma_{bc}]$ $=\eta_{ab}\gamma_{c}-\eta_{ac}\gamma_{b},$ (12c)
$[\gamma_{a},\gamma_{b}\gamma_{5}]$ $=2\eta_{ab}\gamma_{5},$ (12d)
$[\gamma_{a},\gamma_{5}]$ $=2\gamma_{a}\gamma_{5},$ (12e)
$[\sigma_{ab},\sigma_{cd}]$ $=-\eta_{ac}\sigma_{bd}-\eta_{bd}\sigma_{ac}+\eta_{bc}\sigma_{ad}+\eta_{ad}\sigma_{bc},$ (12f)
$[\sigma_{ab},\gamma_{c}\gamma_{5}]$ $=-\eta_{ac}\gamma_{b}\gamma_{5}+\eta_{bc}\gamma_{a}\gamma_{5},$ (12g)
$[\sigma_{ab},\gamma_{5}]$ $=0,$ (12h)
$[\gamma_{a}\gamma_{5},\gamma_{b}\gamma_{5}]$ $=4\sigma_{ab},$ (12i)
$[\gamma_{a}\gamma_{5},\gamma_{5}]$ $=-2\gamma_{a},$ (12j)
$[\gamma_{5},\gamma_{5}]$ $=0.$ (12k)
In bundle constructions over an $n$-dimensional manifold $M$, the underlying object is called the base manifold, while the object defined over it is called the bundle $E$. The base manifold may be regarded as a union of open regions, $M=U\cup V\cup\cdots$, and to each point of $M$ there corresponds a fiber in the bundle, which is generally a $k$-dimensional vector space $V^{k}$. Accordingly, one writes $E=U\times V^{k}$. A map that assigns to each point of $M$ a point in the fiber is called a section, whereas the inverse map is called a projection. Moreover, the group formed by the transformations relating two points in $E$ is called the structure group of $E$ and is denoted by $\mathcal{G}$ [27, 28].
In accordance with the aim of this work, we first summarize the coordinate coframe bundle. On the $(1+3)$-dimensional manifold $M$, through each point pass as many independent curves as the dimension of the manifold, which define a coordinate system $x^{\mu}$, $\mu=0,1,2,3$. The differentials of these coordinates naturally define the objects $dx^{\mu}$, called the coordinate coframe, and each is represented by a point on the fiber. In this context, $dx^{\mu}$ is a section. This bundle is specifically called the coordinate coframe bundle, $CT^{*}(M)=U\times\mathbb{R}^{1,3}$, where $U\subset M$ and $\mathbb{R}^{1,3}$ is a four-dimensional real vector space. On the other hand, at the same point of $M$, one may choose another coordinate system $x^{\mu^{\prime}}$ and construct a new coordinate coframe $dx^{\mu^{\prime}}$. The new and old coordinate coframes correspond to two distinct points on the same fiber of $CT^{*}(M)$. Any two coordinate coframes on the same fiber are related by a linear transformation, $dx^{\mu^{\prime}}=L^{\mu^{\prime}}{}_{\mu}dx^{\mu}$, where the transformation matrix $L^{\mu^{\prime}}{}_{\mu}:=\partial x^{\mu^{\prime}}/\partial x^{\mu}$ is an element of the general linear group $GL(4,\mathbb{R})$. Thus, $GL(4,\mathbb{R})$ is the structure group of $CT^{*}(M)$.
All the above discussion can be repeated for the coordinate frame $\partial/\partial x^{\mu}$, which is dual to the coordinate coframe and satisfies $dx^{\mu}(\partial/\partial x^{\nu})=\delta^{\mu}_{\nu}$. In this setting, $\partial/\partial x^{\nu}$ is the projection of this bundle. Just as we relate two coordinate coframes at the same point of $M$ (that is, two points on the same fiber), we also relate coordinate coframes at two distinct points of $M$ (that is, points on different fibers) by means of the connection of the $CT^{*}(M)$ bundle. Indeed, the connection determines the rule for parallel transport of the coframe. For the bundle construction to be well-defined, the connection 1-forms at two different points of $CT^{*}(M)$ are related through the structure group as $\omega^{\mu^{\prime}}{}_{\nu^{\prime}}=L^{\mu^{\prime}}{}_{\mu}\omega^{\mu}{}_{\nu}L^{\nu}{}_{\nu^{\prime}}+L^{\mu^{\prime}}{}_{\mu}dL^{\mu}{}_{\nu^{\prime}}$, where $L^{\mu}{}_{\mu^{\prime}}:=\partial x^{\mu}/\partial x^{\mu^{\prime}}$ is the inverse of $L^{\mu^{\prime}}{}_{\mu}$, satisfying $L^{\mu}{}_{\mu^{\prime}}L^{\mu^{\prime}}{}_{\nu}=\delta^{\mu}_{\nu}$. This procedure is referred to as covariantization.
Accordingly, the metric measuring the infinitesimal distance between neighboring points on $M$ is expressed in $CT^{*}(M)$ as $g=g_{\mu\nu}(x)dx^{\mu}\otimes dx^{\nu}=g_{\mu^{\prime}\nu^{\prime}}(x^{\prime})dx^{\mu^{\prime}}\otimes dx^{\nu^{\prime}}$, with the components related by $g_{\mu^{\prime}\nu^{\prime}}=L^{\mu}{}_{\mu^{\prime}}g_{\mu\nu}L^{\nu}{}_{\nu^{\prime}}$. Consequently, a $(1,1)$-type tensor-valued exterior form $\mathcal{T}^{\mu}{}_{\nu}$ can be defined on $CT^{*}(M)$, and its covariant exterior derivative is
$D\mathcal{T}^{\mu}{}_{\nu}=d\mathcal{T}^{\mu}{}_{\nu}+\omega^{\mu}{}_{\sigma}\wedge\mathcal{T}^{\sigma}{}_{\nu}-\omega^{\sigma}{}_{\nu}\wedge\mathcal{T}^{\mu}{}_{\sigma}\,$. (13)
Under the transformation rule $dx^{\mu^{\prime}}=L^{\mu^{\prime}}{}_{\mu}dx^{\mu}$, both the tensor components and their covariant exterior derivative transform as
$\mathcal{T}^{\mu^{\prime}}{}_{\nu^{\prime}}=L^{\mu^{\prime}}{}_{\mu}\mathcal{T}^{\mu}{}_{\nu}L^{\nu}{}_{\nu^{\prime}}\qquad\text{and}\qquad D\mathcal{T}^{\mu^{\prime}}{}_{\nu^{\prime}}=L^{\mu^{\prime}}{}_{\mu}\big(D\mathcal{T}^{\mu}{}_{\nu}\big)L^{\nu}{}_{\nu^{\prime}}\,$. (14)
As showed above that structure group of the coordinate frame and coframe is $GL(4,\mathbb{R})$ where this group acts on the frame / coframe canonically. In group algebraic language there exists a standard right-action of $GL(4,\mathbb{R})$ on the fiber, which consists of frames / coframes in this context. Moreoever, one may enrich the structure of the manifold by adding new objects (oftenly come with some constraints). By doing so, according to the new constraints on the manifold, the structure group also receives corresponding restrictions. Generally group reduces to a subgroup, $\mathcal{G^{\prime}}\subset\mathcal{G}$. For instance, in case we equipped the manifold with a general Riemannian metric -which makes the manifold an inner product Riemannian one- it allows us to construct an orthogonal frame bundle over the Riemannian manifold. This process reduces the $GL(4,\mathbb{R})$ to $O(4)\subset GL(4,\mathbb{R})$. Furthermore, by imposing normalization and pseudo-Riemannian conditions as well, we may define orthonormal frame bundle associated with the Lorentz group $SO(1,3)\subset GL(4,\mathbb{R})$ where allows us locally constructing the Lorentz metric by the help of the tetrad (vierbein) in the following way.
Next, by means of the tetrad (vierbein) $h^{a}{}_{\mu}$, we lift $CT^{*}(M)$ to the orthonormal coframe bundle $OT^{*}(M)=U\times\mathbb{R}^{1,3}$ via $e^{a}=h^{a}{}_{\mu}dx^{\mu}$, where $e^{a}\in OT^{*}(M)$, $U\subset M$, and $\mathbb{R}^{1,3}$ is a four-dimensional real vector space. Thus, $e^{a}$ is a section of this bundle. The metric is lifted from $CT^{*}(M)$ to $OT^{*}(M)$ by the inverse tetrad $h^{\mu}{}_{a}$ as $\eta_{ab}=h^{\mu}{}_{a}g_{\mu\nu}h^{\nu}{}_{b}$, where $\eta_{ab}$ are the Minkowski metric components and $h^{a}{}_{\mu}h^{\mu}{}_{b}=\delta^{a}_{b}$. Hence, in $OT^{*}(M)$ the metric reads $g=\eta_{ab}e^{a}\otimes e^{b}$. The connection 1-form is lifted according to $\omega^{a}{}_{b}=h^{a}{}_{\mu}\omega^{\mu}{}_{\nu}h^{\nu}{}_{b}+h^{a}{}_{\mu}dh^{\mu}{}_{b}$. Two orthonormal coframes $e^{a}$ and $e^{a^{\prime}}$ are related by $e^{a^{\prime}}=L^{a^{\prime}}{}_{a}e^{a}$. Requiring invariance of $g=\eta_{ab}e^{a}\otimes e^{b}=\eta_{a^{\prime}b^{\prime}}e^{a^{\prime}}\otimes e^{b^{\prime}}$ yields $\eta_{a^{\prime}b^{\prime}}=L^{a}{}_{a^{\prime}}\eta_{ab}L^{b}{}_{b^{\prime}}$, with $\eta_{a^{\prime}b^{\prime}}=\eta_{ab}=\mathrm{diag}(-1,+1,+1,+1)$. In matrix notation this is $\eta]=[L]^{T}[\eta][L]$, implying that $L^{a^{\prime}}{}_{a}$ forms the Lorentz group $SO(1,3)$. Thus, $SO(1,3)$ is the structure group of $OT^{*}(M)$. Moreover, the $(1,1)$-type tensor-valued exterior form living on $CT^{*}(M)$ and its covariant exterior derivative are lifted to $OT^{*}(M)$ as $\mathcal{T}^{a}{}_{b}=h^{a}{}_{\mu}\mathcal{T}^{\mu}{}_{\nu}h^{\nu}{}_{b}$ and $D\mathcal{T}^{a}{}_{b}=h^{a}{}_{\mu}\big(D\mathcal{T}^{\mu}{}_{\nu}\big)h^{\nu}{}_{b}$, where
$D\mathcal{T}^{a}{}_{b}=d\mathcal{T}^{a}{}_{b}+\omega^{a}{}_{c}\wedge\mathcal{T}^{c}{}_{b}-\omega^{c}{}_{b}\wedge\mathcal{T}^{a}{}_{c}\,$. (15)
Then, under the transformation of the orthonormal coframes $e^{a^{\prime}}=L^{a^{\prime}}{}_{a}e^{a}$, both $\mathcal{T}^{a}{}_{b}$ and its covariant exterior derivative transform in the same manner:
$\mathcal{T}^{a^{\prime}}{}_{b^{\prime}}=L^{a^{\prime}}{}_{a}\mathcal{T}^{a}{}_{b}L^{b}{}_{b^{\prime}}\qquad\text{and}\qquad D\mathcal{T}^{a^{\prime}}{}_{b^{\prime}}=L^{a^{\prime}}{}_{a}\big(D\mathcal{T}^{a}{}_{b}\big)L^{b}{}_{b^{\prime}}\,$. (16)
Finally, by means of the Dirac matrices, we lift the entire structure from the $OT^{*}(M)$ bundle to the spinor bundle $\mathcal{S}(M)=U\times\mathbb{C}^{4}$, where $U\subset M$ and $\mathbb{C}^{4}$ is a four-dimensional complex vector space. The complex vectors defined on this space are called spinors $\psi$. That is, the spinor $\psi$ is a four-component complex column matrix. Hence, each point on a fiber of $\mathcal{S}(M)$ represents a spinor. In other words, $\psi$ is a section of this bundle, and $\overline{\psi}:=\psi^{\dagger}\gamma_{0}$ is its projection, where the dagger symbol denotes Hermitian conjugation. An orthonormal coframe at a point of the $OT^{*}(M)$ bundle can be lifted to $\mathcal{S}(M)$ via the assignment $\gamma=\gamma_{a}e^{a}$. However, lifting the connection 1-form of the $OT^{*}(M)$ bundle to the spinor bundle is not as simple and canonical as lifting the coframe. For the moment, let us denote the connection 1-form of the spinor bundle by $\Omega$ and refer to it as the spinor connection 1-form. Accordingly, the covariant exterior derivative of the spinor is written as
$D\psi=d\psi+\Omega\psi\,.$ (17)
Two spinors $\psi$ and $\psi^{\prime}$ corresponding to two different points on a fiber of $\mathcal{S}(M)$ are related linearly by $\psi^{\prime}=S\psi$, where $S$ is a four-component complex transformation matrix. This matrix $S$ is an element of the structure group of the spinor bundle, $S\in\mathcal{G}$. According to the covariantization rule, in order for $D\psi^{\prime}=S(D\psi)$ to hold under the transformation rule $\psi^{\prime}=S\psi$, the spinor connection 1-form must transform as
$\Omega^{\prime}=S\Omega S^{-1}+SdS^{-1}\,.$ (18)
At the same time, the transformation rule between orthonormal coframes in the spinor bundle can be written as $\gamma^{\prime}=S\gamma S^{-1}$. In this case, two distinct elements $S$ and S-S of the structure group of the spinor bundle map a given orthonormal coframe to the same new coframe, while mapping a spinor to two different spinors. For this reason, we say that the group $\mathcal{G}$ is a double cover of $SO(1,3)$.
For instance, in Einstein-Dirac theory, the lifting is achieved by the assignment $\Omega=\frac{1}{2}\sigma_{ab}\omega^{ab}$. Since in General Relativity the gravitational field is geometrized, $\Omega$ essentially represents the interaction of the spinor with the gravitational field. In this case, $\mathcal{G}=Spin(1,3)$. In Einstein-Dirac-Maxwell theory, one has $\Omega=\frac{1}{2}\sigma_{ab}\omega^{ab}+iqA$, where $q$ is the electric charge of the spinor and $A$ is the Maxwell potential 1-form. In this case, $\mathcal{G}=Spin(1,3)\otimes U(1)$. In Einstein-Cartan-Dirac-Maxwell theory, $\Omega=\frac{1}{2}\sigma_{ab}\omega^{ab}-\frac{1}{2}IT+iqA$, where $T$ is the torsion trace 1-form. In this case, $\mathcal{G}=Spin(1,3)\otimes W(1)\otimes U(1)$, where $W(1)$ denotes the Weyl group [18]. At this stage, it is worth emphasizing that $Spin(1,3)$, $Spin(1,3)\otimes U(1)$, and $Spin(1,3)\otimes W(1)\otimes U(1)$ are all double covers of $SO(1,3)$. Ultimately, the group $\mathcal{G}$ is larger than $SO(1,3)$ yet smaller than $GL(4,\mathbb{R})$. Moreover, depending on the interaction freedom of the spinor, the structure group of the spinor bundle enlarges accordingly.
For the novel covariant exterior derivative of the spinor proposed in (31), the structure group $\mathcal{G}$ is a double cover of the Lorentz group and remains smaller than $GL(4,\mathbb{R})$. The precise structure of this group has not yet been determined; it is reserved as the subject of a separate future investigation.
One may further interpret the proposed novel covariant exterior derivative (31) within the framework of Koszul $\mathcal{G}$-connections. Recall that a Koszul $\mathcal{G}$-connection on a $\mathcal{G}$-structure is characterized by a covariant derivative whose associated connection 1-form takes values in the Lie algebra $\mathfrak{g}$ of the structure group $\mathcal{G}$. In local form, the transformation law (18) ensures precisely that $\Omega$ defines a principal $\mathcal{G}$-connection and hence a Koszul $\mathcal{G}$-connection on the associated vector bundle. In this perspective, the novel covariant exterior derivative (31) may be viewed as defining a new class of Koszul $\mathcal{G}$-connections on the spinor bundle $\mathcal{S}(M)$. In particular, if the corresponding connection 1-form $\Omega$ is $\mathfrak{g}$-valued for a structure group $\mathcal{G}$ that is a double cover of $SO(1,3)$ yet strictly smaller than $GL(4,\mathbb{R})$, then the resulting geometry corresponds to a reduced $\mathcal{G}$-structure endowed with a Koszul $\mathcal{G}$-connection that is not necessarily the lift of the Levi-Civita connection.
Consequently, our novel derivative may generate a geometrically nontrivial extension of the standard spin connection: it defines a Koszul $\mathcal{G}$-connection compatible with the Clifford module structure, while potentially allowing for generalized torsion, non-metricity, or additional internal gauge components encoded in $\mathfrak{g}$. In this sense, (31) does not merely modify the spinor covariant derivative at the level of field theory, but may correspond to a genuinely new class of reduced $\mathcal{G}$-geometries on $M$, whose precise algebraic and geometric characterization remains to be fully determined.
4 Evolution of Dirac equation
Paul Dirac derived the relativistic wave equation for a free partricle with a mass $m$ and a momentum $\vec{p}=-i\hbar(\partial/\partial x,\partial/\partial y,\partial/\partial z)$ in the Cartesian coordinates in 1928
$i\hbar\frac{\partial\psi}{\partial t}=\frac{\hbar c}{i}\left(\alpha_{1}\frac{\partial\psi}{\partial x}+\alpha_{2}\frac{\partial\psi}{\partial y}+\alpha_{3}\frac{\partial\psi}{\partial z}\right)+\beta mc^{2}\psi$ (19)
where $\hbar=h/2\pi$ is the reduced Planck constant, $c$ is the light velocity in vacuum, $\beta,\alpha_{1},\alpha_{2},\alpha_{3}$ are the Dirac matrices and $\psi$ is the Dirac spinor [1, 2]. This equation could be re-casted in the language of exterior forms as
$i*\gamma\wedge d\psi+i\frac{mc}{\hbar}\psi*1=0$ (20)
where * (asterix) is the Hodge map and $\gamma=\gamma_{a}e^{a}$ is $cl(1,3)$-valued 1-form. The appearance of the imaginary unit $i$ is dependent on the choice of gamma matrices and the signature of the metric. The coefficient $i$ in both the kinetic and mass terms should be in our conventions for the Dirac Lagrangian to be hermitian. The hermiticity of the Dirac Lagrangian leads to a charge current which admits the usual probabilistic interpretation [31].
In addition to declaring the “direct Dirac equation”, Lagrangian formulation of the theory in the field theoretical approach is also a very powerful, effective and widely accepted method. Thus, we will name the Dirac equation obtained by variation of the Lagrangian as the “variational Dirac equation” in order to clarify and solidify our arguments in the following sections. Consequently, we write down the Dirac Lagrangian 4-form from which the direct Dirac equation given by Eq.(20) is derived by varying with respect to $\overline{\psi}$ in natural units $\hbar=c=1$
$L_{D}=\frac{i}{2}\left[\overline{\psi}*\gamma\wedge d\psi+d\overline{\psi}\wedge*\gamma\psi\right]+im\overline{\psi}\psi*1$ (21)
where $\overline{\psi}:=\psi^{\dagger}\gamma_{0}$ is the Dirac adjoint spinor. As explained in the previous paragraph, the Dirac Lagrangian is Hermitian. The second term in this Lagrangian can be written as
$d\overline{\psi}\wedge*\gamma\psi=\overline{\psi}*\gamma\wedge d\psi-\overline{\psi}(d*\gamma)\psi+d(\overline{\psi}*\gamma\psi).$ (22)
Since $d*\gamma=0$ in the Cartesian coordinates and the exact derivative does not contribute to the variational field equations, $\overline{\psi}$-variation of the first and second terms yield the same result $\delta\overline{\psi}*\gamma\wedge d\psi$, thus $\overline{\psi}$-variation of the Dirac Lagrangian gives rise to the variational Dirac equation which is the same as the direct Dirac equation given by Eq.(20).
When the spin-1/2 fermion with mass $m$ has also an electric charge $q$ and also interacts with an external electromagnetic field, one can write down the direct Dirac equation by applying the minimal coupling recipe which means replacement of the exterior derivative $d$ with the covariant exterior derivative $\mathcal{D}$,
$i*\gamma\wedge\mathcal{D}\psi+im\psi*1=0\quad\text{where}\quad\mathcal{D}\psi=d\psi+iqA\psi$ (23)
and $A$ electromagnetic potential 1-form. By substituting the minimal coupling paradigm into the Dirac Lagrangian we can write down the following
$L_{DM}=\frac{i}{2}\left[\overline{\psi}*\gamma\wedge\mathcal{D}\psi+\mathcal{D}\overline{\psi}\wedge*\gamma\psi\right]+im\overline{\psi}\psi*1$ (24)
where $\mathcal{D}\overline{\psi}:=\left(\mathcal{D}\psi\right)^{\dagger}\gamma_{0}=d\overline{\psi}-iqA\overline{\psi}$. The variational Dirac equation is obtained by varying this Lagrangian with respect to $\overline{\psi}$. Then, it is seen that the direct Dirac equation and the variational Dirac equation are the same as expected and wished.
Similarly, for a spin-1/2 fermion with mass $m$ interacting with a gravitational field represented by Einstein’s theory of general relativity formulated in the Riemannian spacetime, one can write down the direct Dirac equation via the minimal coupling recipe, i.e., $d\to\widetilde{D}$,
$i*\gamma\wedge\widetilde{D}\psi+im\psi*1=0\quad\text{where}\quad\widetilde{D}\psi=d\psi+\frac{1}{2}\widetilde{\omega}^{ab}\sigma_{ab}\psi.$ (25)
Again, the corresponding Lagrangian is casted readily
$\widetilde{L}_{DE}=\frac{i}{2}\left[\overline{\psi}*\gamma\wedge\widetilde{D}\psi+\widetilde{D}\overline{\psi}\wedge*\gamma\psi\right]+im\overline{\psi}\psi*1$ (26)
where $\widetilde{D}\overline{\psi}:=\left(\widetilde{D}\psi\right)^{\dagger}\gamma_{0}=d\overline{\psi}-\frac{1}{2}\widetilde{\omega}^{ab}\overline{\psi}\sigma_{ab}$. The second term in this Lagrangian can be written as
$\widetilde{D}\overline{\psi}\wedge*\gamma\psi=\overline{\psi}*\gamma\wedge\widetilde{D}\psi-\overline{\psi}(D*e^{a}-\widetilde{\omega}^{(ab)}\wedge*e_{b})\gamma_{a}\psi+d(\overline{\psi}*\gamma\wedge\psi)\;.$ (27)
Since $\widetilde{D}*e^{a}=0$ and $\widetilde{\omega}^{(ab)}=0$ in the Riemannian geometry, the variational Dirac equation is obtained easily from $\overline{\psi}$-variation of the Dirac Lagrangian. Thus, it is seen plausibly that the direct Dirac equation and the variational Dirac equation are the same as hoped and wished.
By following the same strategy in the literature, the direct Dirac equation in the Riemann-Cartan spacetime is written down explicitly
$i*\gamma\wedge\widehat{D}\psi+im\psi*1=0\quad\text{where}\quad\widehat{D}\psi=d\psi+\frac{1}{2}\widehat{\omega}^{ab}\sigma_{ab}\psi$ (28)
and $\widehat{\omega}^{ab}=\widetilde{\omega}^{ab}+K^{ab}$. Then, the appropriate Dirac Lagrangian [10, 14] turns out to be
$\widehat{L}_{DRC}=\frac{i}{2}\left[\overline{\psi}*\gamma\wedge\widehat{D}\psi+\widehat{D}\overline{\psi}\wedge*\gamma\psi\right]+im\overline{\psi}\psi*1$ (29)
where $\widehat{D}\overline{\psi}:=\left(\widehat{D}\psi\right)^{\dagger}\gamma_{0}=d\overline{\psi}-\frac{1}{2}\widehat{\omega}^{ab}\overline{\psi}\sigma_{ab}$. However, the variational Dirac equation derived by varying $\widehat{L}_{DRC}$ with respect to $\overline{\psi}$
$i*\gamma\wedge\left(\widehat{D}\psi-\frac{1}{2}T\psi\right)+im\psi*1=0$ (30)
is not surprisingly the same as the direct Dirac equation given by Eq.(28) for the Riemann-Cartan spacetime. This extraordinary result, which does not manifest itself in Riemannian spacetime, has been studied extensively in various studies with different motivations and names [14, 15, 16, 17, 18] and the references therein. However, at the same time, this distinction is not scoped with Riemann-Cartan spacetime solely, but a general consequence of any metric-affine spacetime [20, 21, 22, 23, 24, 25] and the references therein. We call this discrepancy the inconsistency problem of the Dirac equation in non-Riemannian geometries. One remedy about it was discussed in three-dimensional Riemann-Cartan spacetime in Ref. [18] by extending the covariance group from the Lorentz group to the Lorentz-Weyl group. Although $cl(1,2)$ Clifford algebra has the bases $\{I,\gamma_{a},\sigma_{ab}\}$ in three dimensions, that discussion does not include the term $\gamma_{a}$ in the spinor connection. In addition, non-metricity is never even mentioned in the discussion. Thus, we will handle this inconsistency problem in as wide a perspective as possible in the following section by including all bases of $cl(1,3)$ Clifford algebra in the spinor connection in four dimensions.
5 Our resolution to the inconsistency problem
In this paper, we discuss this inconsistency problem and propose a solution in a novel way in four dimensional metric-affine spacetimes from the most general perspective. We argue that the covariant exterior derivative of the spinor must include not only $\sigma_{ab}$, but also all the bases of $cl(1,3)$ and correspondingly all the possible ingredients of the geometry. Since the dynamics of a metric-affine geometry are characterized only by $e^{a}$ and $\omega^{a}{}_{b}$ in the orthonormal frames, by saying all the possible ingredients of the geometry we mean both. Therefore, first time we suggest
$\mathbb{D}\psi$ $=(d+\Omega)\psi\qquad\text{where}$ (31a)
$\Omega=\frac{1}{2}\sigma_{ab}\omega^{ab}+(a_{1}I+a_{2}\gamma_{5})Q$ $+(a_{3}I+a_{4}\gamma_{5})P+(b_{1}I+b_{2}\gamma_{5})T+(b_{3}I-b_{4}\gamma_{5})\gamma$ (31b)
where $\Omega$ is the spinor connection 1-form, and $a_{i}$ and $b_{i}$, $i=1,2,3,4$, are arbitrary coupling complex constants. Its Dirac adjoint is obtained by usage of the identities given by Eq.(3) as
$\mathbb{D}\overline{\psi}$ $:=\left(\mathbb{D}\psi\right)^{\dagger}\gamma_{0}=d\overline{\psi}-\overline{\psi}(\gamma_{0}\Omega^{\dagger}\gamma_{0})$ (32)
$=d\overline{\psi}+\overline{\psi}\Big[-\frac{1}{2}\sigma_{ab}\omega^{ab}+(a_{1}^{\star}I+a_{2}^{\star}\gamma_{5})Q+(a_{3}^{\star}I+a_{4}^{\star}\gamma_{5})P$
$\qquad\qquad\qquad+(b_{1}^{\star}I+b_{2}^{\star}\gamma_{5})T-(b_{3}^{\star}I+b_{4}^{\star}\gamma_{5})\gamma\Big]$ (33)
where (star) denotes the complex conjugate. Accordingly, the direct Dirac equation for the most general non-Riemannian spacetime is written down
$i*\gamma\wedge\mathbb{D}\psi+im\psi*1=0\;.$ (34)
Then, the corresponding Dirac Lagrangian becomes
${L}_{\mathbb{D}}=\frac{i}{2}\left[\overline{\psi}*\gamma\wedge\mathbb{D}\psi+\mathbb{D}\overline{\psi}\wedge*\gamma\psi\right]+im\overline{\psi}\psi*1.$ (35)
Now, we compute the variational Dirac equation by varying the Lagrangian $L_{\mathbb{D}}$ with respect to $\overline{\psi}$. For this purpose, we rewrite the second term in the Lagrangian as
$\mathbb{D}\overline{\psi}\wedge*\gamma\psi$ $=\overline{\psi}\wedge*\gamma\wedge\Big\{d\psi+\frac{1}{2}\sigma_{ab}\omega^{ab}\psi+(-a_{1}^{\star}I+a_{2}^{\star}\gamma_{5})Q\psi$
$\quad+(-a_{3}^{\star}I+a_{4}^{\star}\gamma_{5})P\psi+(-b_{1}^{\star}I+b_{2}^{\star}\gamma_{5})T\psi+(b_{3}^{\star}I-b_{4}^{\star}\gamma_{5})\gamma\psi\Big\}$
$\quad-\overline{\psi}(D*e^{a}-\omega^{ab}\wedge*e_{b})\gamma_{a}\psi+d(\overline{\psi}*\gamma\wedge\psi).$ (36)
In this step the commutator relations given by Eq.(12) have been used repeatedly. Besides, we insert the following result derived from the identity Eq.(4a) into the above
$D*e^{a}-\omega^{(ab)}\wedge*e_{b}=*e_{a}\wedge(Q+T-P).$ (37)
Then we arrive at
$\mathbb{D}\overline{\psi}\wedge*\gamma\psi$ $=\overline{\psi}\wedge*\gamma\wedge\Big\{d+\frac{1}{2}\sigma_{ab}\omega^{ab}+[-(a_{1}^{\star}+1)I+a_{2}^{\star}\gamma_{5}]Q+[-(a_{3}^{\star}-1)I$
$\quad+a_{4}^{\star}\gamma_{5}]P+[-(b_{1}^{\star}+1)I+b_{2}^{\star}\gamma_{5}]T+(b_{3}^{\star}I-b_{4}^{\star}\gamma_{5})\gamma\Big\}\psi+d(\overline{\psi}*\gamma\wedge\psi).$
Meanwhile, we rewrite expicitly the first term in the Lagrangian given by Eq.(35) as
$\overline{\psi}\wedge*\gamma\wedge\mathbb{D}\psi$ $=\overline{\psi}\wedge*\gamma\wedge\Big\{d+\frac{1}{2}\sigma_{ab}\omega^{ab}+(a_{1}I+a_{2}\gamma_{5})Q+(a_{3}I+a_{4}\gamma_{5})P$
$\qquad+(b_{1}I+b_{2}\gamma_{5})T+(b_{3}I-b_{4}\gamma_{5})\gamma\Big\}\psi.$ (38)
Thus $\overline{\psi}$-variation of the Dirac Lagrangian yields the variational Dirac equation
$i*\gamma\wedge\Big\{d+\frac{1}{2}\sigma_{ab}\omega^{ab}+\frac{1}{2}[(a_{1}-a_{1}^{\star}-1)I+(a_{2}+a_{2}^{\star})\gamma_{5}]Q$
$+\frac{1}{2}[(a_{3}-a_{3}^{\star}+1)I+(a_{4}+a_{4}^{\star})\gamma_{5}]P+\frac{1}{2}[(b_{1}-b_{1}^{\star}-1)I+(b_{2}+b_{2}^{\star})\gamma_{5}]T$
$+\frac{1}{2}[(b_{3}+b_{3}^{\star})I+-(b_{4}+b_{4}^{\star})\gamma_{5}]\gamma\Big\}\psi+im\psi*1$ $=0\;.$ (39)
In order for this variational Dirac equation to be the same as the direct Dirac equation given by Eq.(34) the coupling constants must satisfy
$a_{1}-a_{1}^{\star}-1$ $=2a_{1},$ $a_{2}^{\star}$ $=a_{2},$ $a_{3}-a_{3}^{\star}+1$ $=2a_{3},$ $a_{4}^{\star}$ $=a_{4},$ (40a)
$b_{1}-b_{1}^{\star}-1$ $=2b_{1},$ $b_{2}^{\star}$ $=b_{2},$ $b_{3}^{\star}$ $=b_{3},$ $b_{4}^{\star}$ $=b_{4}.$ (40b)
Under these choices the coupling coefficients could be written in the following form
$a_{1}$ $=-\frac{1}{2}+iA_{1},$ $a_{2}$ $=A_{2},$ $a_{3}$ $=+\frac{1}{2}+iA_{3},$ $a_{4}$ $=A_{4},$ (41a)
$b_{1}$ $=-\frac{1}{2}+iB_{1},$ $b_{2}$ $=B_{2},$ $b_{3}$ $=B_{3},$ $b_{4}$ $=B_{4},$ (41b)
where $A_{1},A_{2},A_{3},A_{4},B_{1},B_{2},B_{3},B_{4}\in\mathbb{R}$.
All arguments and motivations have been on mathematical bases until now. Nevertheless, with the aim of making some physical interpretations, we rewrite the Dirac Lagrangian given by Eq.(35) in the form below
$L_{D}$ $=\mathcal{H}er\left[i\overline{\psi}*\gamma\wedge\mathbb{D}\psi+im\overline{\psi}\psi*1\right].$ (42)
Then we separate the coframe terms and the connection terms in the definition $\mathbb{D}\psi$ given by Eq.(31) as follows
$\mathbb{D}\psi$ $=\widehat{\mathbb{D}}\psi+(b_{3}\gamma_{a}+b_{4}\gamma_{a}\gamma_{5})e^{a}\psi$ (43)
where
$\widehat{\mathbb{D}}\psi:=\left\{d+\frac{1}{2}\sigma_{ab}\omega^{[ab]}+(a_{1}I+a_{2}\gamma_{5})Q+(a_{3}I+a_{4}\gamma_{5})P+(b_{1}I+b_{2}\gamma_{5})T\right\}\psi.$ (44)
Thus, the first term in Eq.(42) becomes
$i\overline{\psi}*\gamma\wedge\mathbb{D}\psi=i\overline{\psi}*\gamma\wedge\widehat{\mathbb{D}}\psi-4i\overline{\psi}(b_{3}+b_{4}\gamma_{5})\psi*1$ (45)
where we used the identity $e^{b}\wedge e^{a}=-\eta^{ab}*1$. We now substitute this result back into Eq.(42):
$L_{D}=\mathcal{H}er\left[i\overline{\psi}*\gamma\wedge\widehat{\mathbb{D}}\psi+i\overline{\psi}(m-4b_{3}-4b_{4}\gamma_{5})\psi*1\right]$ (46)
From the last term, we deduce that the coefficients $b_{3}$ and $b_{4}$ shift the spinor mass. In other words, even if the spinor is massless at the outset, the geometry induces an effective mass. More interestingly, due to $\gamma_{5}$ matrix, the coefficient $b_{4}$ implies that the left-handed and right-handed components of the spinor acquire different masses. A similar conclusion is reached by a totally different perspective in [32]. In addition, there are works arguing for parallel results in the literature. For example, it is shown in [33] that the Standard Model predicts charge-changing weak interactions for right-handed fermions that can be larger than those for left-handed fermions if the mass is sufficiently large, as is the case for the top quark. Similarly, the calculations based on the Standard Model in [34] show that the weak interaction mass of left-handed polarized fermions is always greater than that of right-handed ones in flight with the same speed in any inertial frame. On the experimental side, the authors of [35, 36, 37] present, respectively, the first and then precise and finally high-precision measurements of the left-right cross section asymmetry for $Z$ boson production by positron–electron collisions. Thus, hypothetically it might be concluded that the left-right asymmetry encountered in those measurements is due to geometry (or gravity).
6 Summary and conclusion
We present a unified treatment for the Dirac equation in Minkowski, Riemannian and Riemann–Cartan spacetimes in terms of exterior algebra. The minimal coupling recipe, $d\to D$, central to this study, is performed throughout the formulation. Two distinct procedures are used to derive the Dirac equation: i. The partial derivative is replaced by the covariant derivative directly at the equation level. ii. The variational machinery is operated after replacement $d\to D$ at the Lagrangian level. An a priori expectation of formal equivalence between the two procedures is challenged by the emergence of a non-trivial inconsistency in the Riemann-Cartan spacetime. We propose a new generalized covariant exterior derivative of the spinor, given by Eq.(31), as a solution to this problem. It contains eight arbitrary coupling complex constants. We think that the full spin connection 1-form must include all bases of the Clifford algebra $cl(1,3)$. Accordingly, all possible and suitable dynamical metric-affine geometric objects such as torsion, non-metricity, orthonormal coframe are determined. We arrive at the result that imposing specific constraints on the coupling constants, given by Eq.(41), resolves the inconsistency. A short analysis on the generalized Dirac Lagrangian is performed in order to gain some physical insights. It is found that two of the novel terms introduced in the full spin connection may cause a mass gain or a shift to the spinor, which distinguishes its handedness. Further investigations on physical insights apart from our newly preprint [38] and the structure of the covariance group on the proposed generalized Dirac equation are being studied through our ongoing project.
Acknowledgements: This work has been supported by TÜBİTAK (The Scientific and Technological Research Council of Türkiye) via the project 124F325. MA thanks Tomi Koivisto for fruitful discussions at the workshop of teleparallel universes in Salamanca, Spain, 26-28 November 2018. The authors sincerely thank the anonymous referees for their thorough and attentive review of the manuscript, and for their insightful suggestions and constructive criticisms, which have significantly contributed to improving the clarity and overall quality of the paper.
Data Availability Statement: Data sharing not applicable to this article as no datasets were generated or analyzed during the current theoretical research.
Conflict of Interest: The authors declare no conflict of interest.